# Learning Diffusion Policies for Robotic Manipulation of Timber Joinery under Fabrication Uncertainty

### v1

@@ 2 Methods
This section outlines the experimental setup, data collection process, policy learning approach, and evaluation procedure used in our robotic manipulation system for timber joinery. We implement diffusion policy learning via behavioral cloning using robot pose and F/T data collected from teleoperated demonstrations.
@@ 2.1 Experimental setup
Our experimental setup consists of two primary systems for conducting this research. The first is a multi-robot setup equipped with customized end effectors for executing the timber assembly task. The second is a teleoperation system for manually operating the robotic setup and collecting data for demonstrations. This subsection details the specific hardware used for our experiments and their software interfaces.
The multi-robot setup consists of two 6-axis industrial robotic arms [^1] ABB IRB 4600 ABB Group , each with a 40 kg payload and a 2.55 m reach mounted on linear tracks (Fig. 2). Both robotic arms utilize custom pneumatically-driven end effectors designed for gripping long timber elements with varying cross-sectional profiles, as typically encountered in timber frame construction. One arm grasps the timber mortise in place and can systematically simulate construction deviations by adding positional offsets. The other arm grasps the timber tenon and is the primary agent for manipulation and completing the assembly task.
@@
Because the industrial robotic arms do not provide native access to joint torque values, we developed the end effector on the tenon-grasping arm to include a 6-axis F/T sensor [^2] ATI Delta IP60, with SI-330-30 calibration ATI Industrial Automation to enable F/T data collection (Fig. 3). For safety, the end effector is also equipped with a pneumatic anti-collision sensor [^3] Schunk OPR 081-P00 Schunk , which both provides passive mechanical compliance and triggers a protective stop on the robotic arm under excessive force, safeguarding the robotic setup and F/T sensor. Motion control and pose feedback for the robotic arms utilize ABB’s Externally Guided Motion (EGM) feature ABB Robotics [2017], a UDP-based real-time control interface. F/T data is transferred over EtherCAT to a central Programmable Logic Controller (PLC) [^4] Beckhoff CX2062 Beckoff . All data communications are centralized on a desktop computer employing ROS [^5] ROS 2 Jazzy Jalisco Open Robotics [2023] interfaces and Python wrappers.
@@
Figure 3: Gripper for holding the tenon piece equipped with anti-collision and F/T sensors.
The teleoperation system (Fig. 4) consists of a virtual reality (VR) system [^6] HTC VIVE Pro 2 HTC Corporation , and we developed a unilateral pose-to-pose transformation from the VR controller to the robotic arm’s end effector. This transformation is essential for aligning the coordinate frames of the VR system and the robotic setup, and for compensating for the human demonstrator’s orientation relative to the robot. For instance, when the demonstrator is facing the robotic arm, pushing the controller forward should consistently move the end effector away from the demonstrator, regardless of their position. In addition, this transformation includes an adjustable scaling parameter that maps the demonstrator’s control motions to proportionally smaller movements on the robot, enhancing the demonstrator’s fine motor control capabilities during high-precision tasks such as timber joint assembly. During teleoperation, robot motion and data collection (see Section 2.2) are enabled by the hand controller’s trigger. Similar to clicking and dragging a mouse, holding the trigger moves the robot in the desired direction, and releasing it allows the controller to be repositioned without affecting the robot or recording idle data. This allows for long motions to be performed in smaller, controlled segments. The VR controller interface utilizes OpenVR Valve Corporation [2024] in Python and ROS, integrating directly with the multi-robot setup by streaming mapped poses to EGM for fluid, real-time teleoperation.
@@
2.2 Data collection
To train our diffusion policies, we created datasets consisting of expert demonstrations. An overview of the data collection workflow is illustrated in Fig. 5. During data collection, a human demonstrator utilized the teleoperation system to manipulate a vertical tenon into the mortise. The demonstrator initiates the task by simultaneously straightening the tenon and executing a downward motion to guide it into the mortise. To improve the robustness of the learned policy under fabrication uncertainties, we dedicated approximately 25% of the demonstrations to error recovery scenarios. In these cases, the demonstrator intentionally initiates contact with the edges of the mortise and performs a sliding maneuver along the surface to realign and insert the tenon, mimicking realistic recovery behaviors in the presence of misalignment. Each collected demonstration consists of timestamped robot poses and F/T data during motion (while the VR controller’s trigger is held).
Figure 5: Data collection and policy rollout workflow.
The raw data for each demonstration is recorded in real-time for each sensor interface: pose data is collected from EGM as a seven-dimensional (7D) vector (3D position in millimeters + 4D quaternion rotation of the tenon gripper end effector) every 12 ms (approximately 83 Hz), while F/T data is collected as a 6D vector (3D force in newtons + 3D torque in newton-meters) at 64 Hz after passing through an infinite impulse response (IIR) low-pass filter to attenuate frequency components above 64 Hz. This filtering step mitigates aliasing and improves signal quality by removing sensor noise and dynamic vibrations introduced by high-speed motion or structural resonances. To temporally align the different sample rates to a common time base, the pose and F/T data are interpolated to 60 Hz; position and F/T data are linearly interpolated, while rotation data uses spherical linear interpolation to preserve smooth interpolation.
Finally, each episode is post-processed to reduce unwanted noise in the demonstrations and remove periods of inactivity or unintentional stillness. Removing idle or near-idle periods is critical as it helps the policy training to focus on active, meaningful interactions rather than noise or moments where no significant movement occurs. We filter the pose and F/T data using a low-pass Butterworth filter Butterworth [1930], with filter parameters selected to reduce unwanted high-frequency noise while preserving the underlying movement patterns (Fig. 6). For the pose data, we used a 4th-order low-pass Butterworth filter with a 1 Hz cutoff, which provides a steeper roll-off and ensures that only low-frequency components characteristic of slower, continuous motion trajectories remain. In contrast, we used a 1st-order low-pass Butterworth filter with a higher 10 Hz cutoff to preserve sharper transitions (e.g., at the moment of contact) while effectively removing high-frequency noise that may come from mechanical oscillations or sensor inaccuracies. Actions are then derived from finite differences between successive pose data points, and any resulting vectors with magnitudes below a set threshold are removed to discard idle or near-idle actions.
@@ 2.2 Data collection
2.3 Policy learning
Our research employs CNN-based diffusion policy learning via behavioral cloning Chi et al. [2024a]. The method is based on DDPMs Sohl-Dickstein et al. [2015], Ho et al. [2020] and DDIMs Song et al. [2022], using DDPM for the forward noising process during training and DDIM for the reverse denoising process and faster inference during rollouts. Our diffusion policies predict a sequence of robot actions (i.e., the robot pose) conditional on the observation horizon of robot pose and F/T data. The foundation for our diffusion models is built upon the work of Bogert et al. van den Bogert et al. [2024].
When loading the dataset for training our diffusion policies, trajectories of 64 points are subsampled from each demonstration. This subsampling normalizes the trajectory length for each demonstration, as well as enables faster training iterations. Next, we converted pose quaternions to a continuous 6D rotation representation (by first converting to rotation matrices then concatenating the first two columns) to improve training stability Zhou et al. [2019], Chi et al. [2024a], resulting in 9D pose vectors. We then applied min-max normalization to scale all features to the range [-1, 1]. Given a pose observation horizon $T_{o}^{p}$, F/T observation horizon $T_{o}^{f}$, and action prediction horizon $T_{p}$, each training sample at timestep $t$ consists of an observation feature $\mathbf{O}_{t}$ concatenating the last $T_{o}^{p}$ and $T_{o}^{f}$ and a corresponding action sequence $\mathbf{A}_{t}$ consisting of the next $T_{p}$ absolute poses.
During the rollouts, $\mathbf{A}_{t}$ is predicted conditioned on $\mathbf{O}_{t}$. The reverse denoising process for action prediction starts from ${A}^{K}$ sampled from Gaussian noise. After iterating $K$ denoising steps, we output a desired noise-free action ${A}^{0}$ at step 0. The action at each reverse step k1k-1 is calculated as Chi et al. [2024a]:
@@ 2.3 Policy learning
where $\epsilon^{k}$ is the sampled random noise with known variance at iteration $k$. The predicted noise $\epsilon_{\theta}(\mathbf{O}_{t},\mathbf{A}^{k}_{t},k)$ is a function of $\mathbf{O}_{t}$, the pose and F/T observation feedback preceding time $t$, $\mathbf{A}^{k}_{t}$ = $\mathbf{A}^{0}+\epsilon^{k}$ with $\mathbf{A}^{0}$ sampled from starting noise-free actions, and iteration step $k$. As implemented in Chi et al. [2024a], we add the sampled random noise $\epsilon^{k}$ to the noise-free actions $\mathbf{A}_{0}$ in $K$ forward steps (original DDPM method).
@@
Policy training was conducted on high-performance computing (HPC) clusters [^7] Princeton Research Computing’s Della cluster, which provide powerful computational resources for efficiently processing large datasets and enabling faster, scalable learning by leveraging parallel computing capabilities. Each policy was trained for a fixed amount of time (1 hour), with each training lasting at least 1,000 epochs. Only the model weights corresponding to the epoch with the lowest validation loss were used to evaluate performance during rollouts. To account for training variance, we trained 4 models for each set of parameters and averaged their success rates when computing their performance (see Section 2.4).
2.4 Evaluation
@@
To evaluate the performance of our trained policies, we perform rollouts on the experimental setup (Section 2.1). The starting position for the tenon piece during all rollouts is fixed and identical to the demonstrations. During policy rollouts, we specify additional parameters for the number of inference steps $K_{\text{inf}}$, DDIM/DDPM interpolation $\eta$ Song et al. [2022], and action execution steps $T_{a}$. The number of inference steps ($K_{\text{inf}}\leq K$) controls the number of steps in the reverse denoising process during policy inference. Decreasing $K_{\text{inf}}$ reduces computation and increases rollout speed, which is important for real-time policy execution. However, smaller $K_{\text{inf}}$ may also reduce the quality of noise sampling or action precision.
@@
The amount of stochasticity during the reverse denoising process is controlled by the interpolation parameter $\eta\in[0,1]$, where $\eta=0$ is the DDIM process and $\eta=1$ is the original DDPM process. This means that when $\eta=0$, the standard deviation of the noise added during sampling is 0, meaning that the reverse process is fully deterministic, producing high-fidelity imitations of demonstrations, and when $\eta=1$, the reverse process is fully stochastic, introducing broader behavioral coverage with exploration and variability.
@@
The number of action execution steps $T_{a}\leq T_{p}$ denotes that of the $T_{p}$ predicted actions at each time step; we execute the first $T_{a}$ actions. Decreasing the value of $T_{a}$ increases the frequency of replanning over a receding horizon, increasing the responsiveness of the system; however, it can also potentially decrease the speed of rollouts by spending more time on inference. For values of $T_{a}>2$, we apply an average pooling with a kernel size of $T_{a}-1$ across each dimension of the actions, smoothing the action sequence.
@@
In contrast to data collection during demonstrations, observations during rollouts are not temporally aligned and instead utilize the latest available pose and F/T data when the observation is retrieved. These observations are normalized using the same normalization parameters as the dataset before being passed to the policy for inference. The output action sequence is similarly denormalized back into the original scale using the same action normalization parameters during training. These predicted actions are then sent through EGM to the robot controller, and we use a similar communication protocol to the data collection to receive pose and F/T feedback (see Fig. 5).
The policy performance is determined by rolling out 5 times for each of the 4 trained models with fixed parameters, to account for stochasticity during test time, resulting in an average success rate (Avg. SR) across 20 combined rollouts. A rollout is recorded as a success if the tenon element fully inserts into the mortise hole without triggering any collision errors (i.e., excessive force at either the collision sensor or the robot joints). A rollout is recorded as a failure if any collision errors are triggered by the industrial robot controller or if the policy does not appear to be making any meaningful progress towards task completion, as determined by the operator.
@@ 3 Experiments
We conducted a two-phase study: in Phase 1 we performed static experiments with a fixed mortise location to establish a general performance evaluation for the diffusion policy in our industrial setup; in Phase 2 we performed uncertainty experiments with randomized hole offsets to evaluate the robustness of the diffusion policies.
The dimensions of the mortise and tenon we used for the experiments are illustrated in Fig. 7. The tolerance between the tenon cheeks and the mortise walls was 0.1 mm, and the tolerance between the tenon edges and the mortise walls was about 2 mm. The tight tolerance of the tenon cheeks is what ensures a structural fit for a mortise and tenon joint. This figure also visualizes the boundary of the 5 mm and 10 mm offsets used in the uncertainty experiments (see Section 3.2). The tenon is aligned along the world’s z-axis, while the mortise piece lies horizontally along the x-axis. Each demonstration and rollout begins with the tenon element grasped at a fixed starting pose, positioned approximately 15 mm vertically above the mortise hole with an approximately 6°angular deviation about the y-axis.
@@
Figure 7: Mortise and tenon terminology, dimensions, and reference frame.

### v2

@@ 3 Methods
This section outlines the experimental setup, including the multi-robot workcell, communication stack, and teleoperation system to collect expert demonstrations (Section 3.1), data synchronization and processing methods to prepare the dataset for training (Section 3.2), a detailed overview of the sensory–motor diffusion policy training method (Section 3.3), and the policy inference details to evaluate the trained policies on our experimental setup (Section 3.4).
@@ 3.1 Experimental setup
Our experimental setup consists of two primary systems for conducting this research. The first is a multi-robot setup equipped with customized end effectors for contact-rich manipulation and timber assembly. The second is a teleoperation system for manually operating the robotic setup and collecting data during demonstrations. This subsection details the specific hardware used for our experiments and their software interfaces.
The multi-robot setup consists of two six-axis industrial robotic arms [^1] ABB IRB 4600 ABB Group , each with a 40 kg payload and a 2.55 m reach mounted on linear tracks (Fig. 2). Both robotic arms utilize custom pneumatically-driven end effectors designed for gripping long elements with varying cross-sectional profiles. Because the industrial robotic arms do not provide native access to joint torque values, one of the end effectors includes a six-axis F/T sensor [^2] ATI Delta IP60, with SI-330-30 calibration ATI Industrial Automation to enable F/T data streaming (Fig. 3). For safety, the same end effector is also equipped with a pneumatic anti-collision sensor [^3] Schunk OPR 081-P00 Schunk , which both provides passive mechanical compliance and triggers a protective stop on the robotic arm under excessive force, safeguarding the robotic setup and F/T sensor.
@@
Motion control and pose feedback for the robotic arms utilize ABB’s Externally Guided Motion (EGM) feature ABB Robotics [2017], a UDP-based real-time control interface. F/T data is transferred over EtherCAT to a central Programmable Logic Controller (PLC) [^4] Beckhoff CX2062 Beckoff . All data communications are centralized on a desktop computer connected to the central PLC via a local area network (LAN) employing ROS [^5] ROS 2 Jazzy Jalisco Open Robotics [2023] interfaces and Python wrappers.
@@
Figure 3: Custom gripper end effector equipped with anti-collision and F/T sensors.
The teleoperation system (Fig. 4) implemented in this study is based on a virtual reality (VR) interface and follows the framework previously introduced in our prior work Ruan et al. [2026]. In this system, six-degree-of-freedom hand poses are streamed from the VR hand controller [^6] HTC VIVE Pro 2 HTC Corporation to the desktop computer via the OpenVR Valve Corporation [2024] application programming interface (API) and mapped in real time to robot end-effector commands streamed over EGM. Specifically, we implement a unilateral pose-to-pose transformation that aligns the VR controller frame with the robot tool center point (TCP) frame. This transformation ensures consistent directional behavior between human input and robot motion (e.g., forward controller motion always results in motion away from the operator, regardless of physical orientation). When the operator presses the controller trigger, the current hand pose and robot TCP pose are recorded as reference frames. While the trigger remains engaged, subsequent controller motions are interpreted as relative displacements with respect to these references and converted into incremental robot motion commands. Translational and rotational components are processed separately and may be independently scaled, enabling fine-grained control during high-precision, contact-rich operations such as tight-tolerance insertion. Additionally, the operator receives haptic feedback through controller vibrations when the measured force from the F/T sensor exceeds a predefined threshold (less than the sensing range maximum value). This cue provides awareness of contact intensity during teleoperation, supporting safer interaction and more controlled force application.
@@
3.2 Data processing
Figure 5: Data collection via teleoperation.
During teleoperation, the raw data for each demonstration is recorded for each sensor interface (Fig. 5): pose data is collected from the robot controller as a seven-dimensional (7D) vector (3D position in millimeters + 4D quaternion rotation of the tenon gripper end effector) every 12 ms (approximately 83 Hz), while F/T data is collected as a 6D vector (3D force in newtons + 3D torque in newton-meters) at 64 Hz after passing through an infinite impulse response (IIR) low-pass filter to attenuate frequency components above 64 Hz. This filtering step mitigates aliasing and improves signal quality by removing sensor noise and dynamic vibrations introduced by high-speed motion or structural resonances. To temporally align the different sample rates to a common time grid, the pose and F/T data are interpolated to 60 Hz; position and F/T data are linearly interpolated, while rotation data uses spherical linear interpolation to preserve smooth interpolation.
Finally, each episode is post-processed to remove periods of inactivity or unintentional stillness and reduce unwanted noise in the demonstrations. Removing idle or near-idle periods is critical as it helps the policy training to focus on active, meaningful interactions rather than noise or moments where no significant movement occurs. We filter the pose and F/T data using a low-pass Butterworth filter Butterworth [1930], with filter parameters selected to reduce unwanted high-frequency noise while preserving the underlying movement patterns (Fig. 6). For the pose data, we used a 4th-order low-pass Butterworth filter with a 1 Hz cutoff, which provides a steeper roll-off and ensures that only low-frequency components characteristic of slower, continuous motion trajectories remain. In contrast, we used a 1st-order low-pass Butterworth filter with a higher 10 Hz cutoff to preserve sharper transitions (e.g., at the moment of contact) while effectively removing high-frequency noise that may come from mechanical oscillations or sensor inaccuracies.
@@ 3.2 Data processing
3.3 Policy training
We train a CNN-based sensory–motor diffusion policy via action diffusion Chi et al. [2024a]. The policy training uses a standard DDPM-style training Ho et al. [2020]. During inference, we adopt a DDIM-style sampling procedure Song et al. [2022], which enables accelerated trajectory generation using fewer denoising steps for efficient action prediction during robot execution. The policy predicts a sequence of robot actions (i.e., the robot TCP pose) conditional on an observation horizon of robot pose and F/T data. The foundation for our diffusion policy is built upon the work of Bogert et al. van den Bogert et al. [2024].
When loading the dataset for training the diffusion policies, trajectories of 64 points are subsampled from each demonstration. This subsampling normalizes the trajectory length for each demonstration, as well as enables faster training iterations. We also convert pose quaternions to a continuous 6D rotation representation (by first converting to rotation matrices, then concatenating the first two columns) to improve training stability Zhou et al. [2019], Chi et al. [2024a], resulting in 9D pose vectors. We then apply min-max normalization to scale all features to the range [-1, 1]. Below, we formalize the diffusion policy training on the subsampled and normalized dataset.
Given a pose observation horizon $T_{o}^{p}$, F/T observation horizon $T_{o}^{f}$, and action prediction horizon $T_{p}$, each training sample at timestep $t$ consists of an observation feature $\mathbf{O}_{t}$ concatenating the last $T_{o}^{p}$ and $T_{o}^{f}$ and a corresponding action sequence $\mathbf{A}_{t}$ consisting of the next $T_{p}$ absolute poses. During inference, $\mathbf{A}_{t}$ is predicted conditioned on $\mathbf{O}_{t}$. The reverse denoising process for action prediction starts from ${A}^{K}$ sampled from Gaussian noise. After iterating $K$ denoising steps, we output a desired noise-free action ${A}^{0}$ at step 0. The action at each reverse step k1k-1 is calculated as Chi et al. [2024a]:
@@ 3.3 Policy training
where $\epsilon^{k}$ is the sampled random noise with known variance at iteration $k$. The predicted noise $\epsilon_{\theta}(\mathbf{O}_{t},\mathbf{A}^{k}_{t},k)$ is a function of $\mathbf{O}_{t}$, the pose and F/T observation feedback preceding time $t$, $\mathbf{A}^{k}_{t}$ = $\mathbf{A}^{0}+\epsilon^{k}$ with $\mathbf{A}^{0}$ sampled from starting noise-free actions, and iteration step $k$. As implemented in Chi et al. [2024a], the sampled random noise $\epsilon^{k}$ is added to the noise-free actions $\mathbf{A}_{0}$ in $K$ forward steps (original DDPM method).
@@
Policy training was conducted on high-performance computing (HPC) clusters [^7] Princeton Research Computing’s Della cluster, which provide powerful computational resources for efficiently processing large datasets and enabling faster, scalable learning by leveraging parallel computing capabilities. Each policy was trained for a fixed amount of time (1 hour), with each training using a batch size of 128 and lasting at least 1,000 epochs. Only the model weights corresponding to the epoch with the lowest validation loss were used to evaluate performance during rollouts.
3.4 Policy inference
@@
To evaluate the performance of our trained policies, we perform rollouts on the experimental setup (Section 3.1). During policy inference, we specify additional parameters for the number of inference steps $K_{\text{inf}}$, DDIM/DDPM interpolation $\eta$ Song et al. [2022], and action execution steps $T_{a}$. The number of inference steps ($K_{\text{inf}}\leq K$) controls the number of steps in the reverse denoising process during policy inference. Decreasing $K_{\text{inf}}$ reduces computation and increases rollout speed, which is important for real-time policy execution. However, smaller $K_{\text{inf}}$ may also reduce the quality of noise sampling or action precision.
@@
The amount of stochasticity during the reverse denoising process is controlled by the interpolation parameter $\eta\in[0,1]$, where $\eta=0$ is the DDIM process and $\eta=1$ is the original DDPM process. This means that when $\eta=0$, the standard deviation of the noise added during sampling is 0, making the reverse process fully deterministic and producing high-fidelity imitations of demonstrations. When $\eta=1$, the reverse process is fully stochastic, introducing broader behavioral coverage through exploration and variability.
@@
The number of action execution steps $T_{a}\leq T_{p}$ denotes that of the $T_{p}$ predicted actions at each time step; we execute the first $T_{a}$ actions. Decreasing the value of $T_{a}$ increases the frequency of replanning over the receding horizon, thereby increasing the system’s responsiveness; however, it may also reduce the speed of rollouts by increasing inference time. For values of $T_{a}>2$, we apply average pooling with kernel size $T_{a}-1$ across each action dimension, thereby smoothing the action sequence.
@@
In contrast to data collection during demonstrations, observations during inference are not temporally aligned, and instead, we use the latest available pose and F/T data when the observation is retrieved. These observations are normalized using the same normalization parameters as the dataset before being passed to the policy for inference. The output action sequence is similarly denormalized back into the original scale using the same action normalization parameters during training.
@@ 4 Experiments
This section describes the experimental methodology and evaluation protocol used in this study. We first introduce the mortise and tenon assembly task and the criteria used to measure policy performance. We then present the data collection procedure used to construct the demonstration datasets. Finally, we outline two experimental phases: Phase 1 focuses on baseline evaluation and parameter initialization under fixed conditions, while Phase 2 fine-tunes the parameters, examines policy robustness under randomized mortise position perturbations, including additional force ablation and demonstration count studies.
4.1 Task Description
The experimental task is the contact-rich robotic assembly of a vertical mortise and tenon joint as illustrated in Fig. 7. Each robot initially grasps a timber element, and we assume that a conventional motion planner is responsible for placing the grasped tenon at a pre-insertion pose. This pre-insertion pose is constant for all demonstrations and rollouts, located near the mortise, positioned approximately 15 mm above the mortise hole, and subject to an initial angular misalignment of approximately 6 degrees about the y-axis (Fig. 7). Uncertainties associated with grasping are outside the scope of this study and are not considered. The task objective is to fully insert the tenon into the mortise using compliant, contact-aware motion. Only the robot with the grasped tenon is controlled while the mortise pose remains fixed throughout the task.
The geometric dimensions of the mortise and tenon used in all experiments are shown in Fig. 7. The clearance between the tenon cheeks and the mortise walls is 0.1 mm, representing a tight structural fit typical of traditional timber joinery, while the clearance between the tenon edges and the mortise walls is approximately 2 mm. The tight tolerance of the tenon cheeks governs load transfer and frictional contact during insertion and is the primary source of contact sensitivity in the task. Figure 7 also visualizes the boundary corresponding to the 5 mm and 10 mm offsets used in the Phase 2 experiments (Section 4.4). These offsets mimic realistic deviations arising from positioning errors, tolerance accumulation, and material imperfections commonly encountered in construction-scale assembly.
@@ 4.1 Task Description
Figure 7: Mortise and tenon assembly task (left), and terminology, dimensions, and reference frame (right).
For each experimental configuration, the policy is trained 4 times with randomized seeds, and policy performance is determined by rolling out 5 times for each trained model, to account for stochasticity during test time, yielding an average success rate (Avg. SR) across 20 combined rollouts. A rollout is recorded as a success if the tenon element fully inserts into the mortise hole without triggering any collision errors (i.e., excessive force at either the collision sensor or the robot joints). A rollout is recorded as a failure if the industrial robot controller triggers any collision errors or the policy does not appear to make meaningful progress toward task completion, as determined by the operator (see examples of successful and failed rollouts in Fig. 9).
4.2 Data collection
To train the diffusion policies, we created two datasets consisting of expert demonstrations. During data collection, a human demonstrator utilized the teleoperation system (Section 3.1) to complete the vertical mortise and tenon assembly task. The first dataset consists of 100 demonstrations with fixed starting positions for the mortise and tenon. This dataset is used by the Phase 1 experiments to initialize parameter tuning (Section 4.3).
The second dataset consists of 400 demonstrations with a fixed tenon starting position but a randomized mortise position, simulating fabrication uncertainty. The mortise position was perturbed by applying a planar translational offset sampled in polar coordinates, with direction $\theta\sim U(0,2\pi)$ and magnitude $r\sim U(0,10\text{~mm})$, and converted to Cartesian offsets $\left(\Delta x,\Delta y\right)=(r\cos{\theta},r\sin{\theta})$. This sampling scheme was selected to approximate assembly tolerances, placing greater probability mass near small offsets while still demonstrating larger deviations and boundary cases, and remaining simple to implement directly on the industrial robot controller. To improve the robustness of the learned policy under fabrication uncertainties, we dedicated 100 (25%) of the demonstrations to error recovery scenarios Black et al. [2026]. In these cases, the demonstrator intentionally initiates contact with the edges of the mortise and slides along the surface to realign and insert the tenon, mimicking realistic recovery behaviors in the presence of misalignment. This dataset is used in the Phase 2 experiments (Section 4.4).